OpenStudy (anonymous):

(tan^2)/(cscx+1).....simplify this trig function

OpenStudy (danjs):

rewrite those qwith only sin and cos

OpenStudy (anonymous):

(sinx/cosx)/(1/sinx +1)???

OpenStudy (danjs):

yeah that makes it worse

OpenStudy (danjs):

i guess maybe, depends

OpenStudy (xapproachesinfinity):

\[\text{recall that }~~~\tan=\frac{\sin}{\cos} ~~and \csc = \frac{1}{\sin } \]

OpenStudy (danjs):

\[\frac{ \sin^3(x) }{ (\sin(x) + 1)*(\cos^2(x)) }\]

OpenStudy (anonymous):

thats as far as i get

OpenStudy (danjs):

the tangent was squared, you forgot that,

OpenStudy (anonymous):

oh yeah.

OpenStudy (danjs):

i dont think i can remember an identity to clean that up more

OpenStudy (anonymous):

can we factor??

OpenStudy (xapproachesinfinity):

you can still clean it up more

OpenStudy (danjs):

unless you want to just make it look different, not sure

OpenStudy (xapproachesinfinity):

yeah more like it but that is fine as well

OpenStudy (anonymous):

so is that as simplified as it gets??

OpenStudy (xapproachesinfinity):

yeah that's fine

OpenStudy (anonymous):

ok cool thanks!

OpenStudy (anonymous):

Thanks!

OpenStudy (danjs):

some examples